### v2

# Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Bias Detection

Abstract
This work contributes to the expanding research on the applicability of LLMs in social sciences by examining the performance of GPT-3.5 Turbo, GPT-4, and Flan-T5 models in detecting framing bias in news headlines through zero-shot, few-shot, and explainable prompting methods. A key insight from our evaluation is the notable efficacy of explainable prompting in enhancing the reliability of these models, highlighting the importance of explainable settings for social science research on framing bias. GPT-4, in particular, demonstrated enhanced performance in few-shot scenarios when presented with a range of relevant, in-domain examples. FLAN-T5’s poor performance indicates that smaller models may require additional task-specific fine-tuning for identifying framing bias detection. Our study also found that models, particularly GPT-4, often misinterpret emotional language as an indicator of framing bias, underscoring the challenge of distinguishing between reporting genuine emotional expression and intentionally use framing bias in news headlines. We further evaluated the models on two subsets of headlines where the presence or absence of framing bias was either clear-cut or more contested, with the results suggesting that these models’ can be useful in flagging potential annotation inaccuracies within existing or new datasets. Finally, the study evaluates the models in real-world conditions ("in the wild"), moving beyond the initial dataset focused on U.S. Gun Violence, assessing the models’ performance on framed headlines covering a broad range of topics111All data created in this study will be made publicly available..
Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Bias Detection
Valeria Pastorino, Jasivan A. Sivakumar, Nafise Sadat Moosavi Department of Computer Science University of Sheffield United Kingdom {vpastorino1|jasivakumar1|n.s.moosavi}@sheffield.ac.uk
1 Introduction
In today’s digital age, the proliferation of news sources and the rapid dissemination of information highlight the essential need for unbiased reporting. Framing bias, which manipulates news content to potentially shift public perception, presents a formidable obstacle in maintaining a well-informed and unbiased public sphere Binotto and Bruno (2018). This form of bias not only influences how events are portrayed but also impacts public attitudes and policy decisions. For instance, the reporting on a government policy change might be emotionally charged in one headline as “Government’s heartless cutbacks leave thousands without essential services”, whereas another might offer a more neutral perspective: “Government announces reduction in funding for public services”.
The concept of framing has received extensive scrutiny in the social sciences, identified as a technique that selectively highlights certain realities to shape audience perceptions and reactions Goffman (1974); Entman (1993). While its significant effect on public communication and media is well-documented, the analysis of framing has historically been challenged by its fragmented nature and reliance on manual, small-scale studies. This situation underscores the necessity for new methods capable of dissecting the complexities of framing in media comprehensively. Amid this landscape, with an overwhelming array of news sources and headlines, it becomes crucial to determine whether they convey biased messages, intentionally or inadvertently. This determination is not merely academic but essential for preserving the integrity of public discourse.
The introduction of large language models (LLMs) opens up novel pathways for identifying framing bias, especially in news headlines. The accessibility and advanced capabilities of models, particularly those within the GPT series, have garnered attention from professionals across various fields, including media analysis and political science, seeking to navigate the extensive corpus of available information. However, despite their impressive success across many applications, the reliability of these models in executing nuanced tasks like detecting framing bias remains an open question. In this study, we address this gap by providing an in-depth investigation into the strengths and shortcomings of state-of-the-art natural language processing (NLP) models in identifying framing bias. We perform a comprehensive examination of the performance of two top-tier NLP models, GPT-4 and GPT-3.5, alongside the open-source FLAN-T5 model (Wei et al., 2022), which has a reasonable performance across different NLP tasks.
Our findings reveal interesting patterns: while the pre-trained FLAN-T5 struggles with this task, the GPT models show promising results. Our results suggest that prompting the models to explain their decisions results in more reliable performances across different settings. A standout discovery is the enhanced reliability of predictions when models are prompted to explain their reasoning, a technique that also narrows the variability in outcomes across different example sets. To test the real-world applicability of these models, apart from a standard dataset that is annotated by experts, we automatically collect a new evaluation set containing framed headlines from different domains. The GPT models also accurately identifying the majority of framed titles in this dataset.
Our analysis suggests that while the examined GPT models achieve high scores on the examined evaluation sets, there are still areas that they struggle. As an example, GPT-4 often mistook emotional language for framing bias. This insight points to the need for developing more nuanced test sets that can present more complex challenges for evaluation. Additionally, our experiments suggest a potential use case for these models beyond mere bias detection. When GPT models, particularly in a more reliable prompting setting, converge on a prediction that deviates from the established gold standard, it may hint at underlying annotation inaccuracies within the dataset. Our findings suggest that a future direction involves constructing more challenging datasets that mirror the real-world complexities of framing bias, as well as enhancing the performance of smaller and open-source models on this challenging task.
2 Related Work
2.1 Automatic Framing Detection
Various NLP methodologies have been used to automate framing bias detection. Early studies in framing detection primarily utilized Topic Modeling, Structural Topic Modeling, and Hierarchical Topic Modeling to uncover themes and topics within large datasets DiMaggio et al. (2013); Nguyen et al. (2015); Gilardi et al. (2021). These methods, while effective in identifying “what” is being discussed, often fall short in revealing “how” information is framed. Latent Dirichlet Allocation (LDA) Topic Modeling, for instance, served as a starting point for creating lists of frames deductively in tools like the one presented by Bhatia et al. (2021) for computational framing analysis. However, as noted by Ali and Hassan (2022), the emphasis in such approaches remains on detecting topics rather than the nuanced framing of those topics. The focus on topics also derives from the connections between agenda setting and framing strategies in computational social sciences, with studies analysing these two phenomena together Field et al. (2018). Further analyses have included pragmatics cues, examining how specific word choices, like the use of “again” in “Again, Dozens of Refugees Drowned”, subtly influence reader perception Yu (2022). This shift towards granular analysis is complemented by advanced models, including Neural Network and deep learning techniques, which offer refined tools for detecting framing nuances Burscher et al. (2016); Card et al. (2015); Liu et al. (2019); Mendelsohn et al. (2021). Tourni et al. (2021) demonstrated that combining transformers models for processing news headlines with residual network models to process news lead images could lead to accurate framing bias detection. Similarly, Naderi and Hirst (2017) explored the use of various deep neural networks, such as LSTMs, BiLSTMs, and GRUs, for frame prediction at the sentence level using the Media Frame Corpus (MFC) Card et al. (2015). Building on this foundation, Liu et al. (2019) and Akyürek et al. (2020) fine-tuned BERT Devlin et al. (2019) to predict frames in news headlines. Their work resulted in the creation of the Gun Violence Frame Corpus (GVFC), a benchmark dataset for framing analysis which will be further discussed in section 2.1.1. More recently, in a novel approach to automatic framing detection, Lai et al. (2022) introduced an unsupervised learning method leveraging Wikipedia’s category system to identify frames within news articles, using the GVFC to evaluate this method.
2.1.1 Datasets
Media Frame Corpus (MFC)
MFC (Card et al., 2015) is a collection of annotated U.S. newspaper articles on topics like immigration, smoking, and same-sex marriage, analyzed for framing. Utilising the Policy Frames Codebook (PFC) by Boydstun et al. (2014), the MFC adopts 14 frame dimensions such as “security and defense” and “cultural identity” for categorising policy discourse. Despite achieving an inter-coder reliability (ICR) of 0.60, critiques, particularly from Ali and Hassan (2022), argue that the PFC’s broad dimensions conflate topics with frames, potentially missing nuanced strategic framing. Moreover, MFC categorises content into wide-ranging dimensions that might not always precisely capture the specific framing intended by a news headline. This categorisation can make it difficult to directly identify whether and how a headline is framed without a deeper, nuanced analysis.
Gun Violence Frame Corpus (GVFC)
Another significant dataset in the field of framing analysis is the Gun Violence Frame Corpus (GVFC), introduced by Liu et al. (2019). This dataset concentrates on the issue of Gun Violence in the U.S. The creation process began with defining nine distinct “frames” related to the topic, drawing from existing literature and a preliminary data analysis. A specialised codebook was then developed, serving as a training tool for annotators along with annotation guidelines. GVFC is made of 2990 news headlines, with 1,300 headlines specific to the issue of Gun Violence in the United States. All the headlines are coded to have a primary frame, while only 319 have 2 frames. For instance, the headline “It’s Time to Hand the Mic to Gun Owners” is annotated with “Public opinion” as the first frame and “2nd Amendment” as the second frame. Similarly, “Trevor Noah: The Second Amendment Is Not Intended for Black People” is annotated with “2nd Amendment” and “Race/Ethnicity” frames (Liu et al., 2019, p. 507).
Non-English Data
Expanding the scope of framing analysis to non-English content, Akyürek et al. (2020) introduced a multilingual extension of the Gun Violence Frame Corpus, which encompasses news headlines in German, Turkish, and Arabic, focusing on U.S. gun violence. This extension involved training two native speakers per language to annotate headlines—350 in German, 200 in Turkish, and 210 in Arabic.
Piskorski et al. (2023b) presented an annotated dataset made of articles spanning though nine languages: English, French, German, Georgian, Greek, Italian, Polish, Russian, and Spanish. This dataset addresses a variety of topics including the COVID-19 pandemic, abortion-related legislation, migration, Russo-Ukrainian war, and various parliamentary elections. The annotation process made use of the PFC codebook, using the 15 dimensions as frames Piskorski et al. (2023a).
While our study focuses on the English language, exploring framing bias detection in less resourced languages presents a promising future direction. By establishing a foundation in English, we lay the groundwork for future research to investigate the strengths and shortcomings of framing bias detection in languages beyond English.
2.2 Evaluating LLMs in Social Science
Applying Large Language Models to social science tasks, such as evaluating sociability Choi et al. (2023), morality Abdulhai et al. (2023), and controversial issues and bias Sun et al. (2023), has received increasing interest, showcasing a wide range of strengths and limitations unique to each task. This diversity stems from the specific challenges and nuances of social phenomena. Although LLMs excel in generating and understanding human-like text, the complex requirements of social science tasks necessitate a detailed, task-specific examination of their performance and reliability.
In this work, we contribute to the expanding research on the applicability of LLMs in social sciences by specifically investigating their reliability in detecting framing bias. This exploration not only aims to identify the strengths and shortcomings of state-of-the-art models in a crucial area of framing analysis but also to extend our comprehension of how these models can be optimised for the nuanced task of framing detection.
3 Experimental Setup
3.1 Data
For our evaluation, we selected the Gun Violence Framing Corpus (GVFC) dataset Liu et al. (2019), motivated by its comprehensive coverage of U.S. Gun Violence framing through 2900 annotated newspaper headlines as well as the high ICR met in the annotation process.222In our experiments, we have excluded the headlines that are not relevant to Gun Violence. Furthermore, we have excluded 22 relevant headlines in order to use them for Few-Shot in domain prompting, leaving 2594 relevant headlines for our analysis. The dataset’s annotations identify whether each headline reflects any of nine critical aspects of gun violence framing: gun rights, gun control, politics, mental health, public/school safety, race/ethnicity, public opinion, social/cultural issues, and economic consequences. In our study, headlines tagged with any of these framing aspects were categorised as framed, while all others were classified as not framed.
3.2 Models
To assess the capabilities of state-of-the-art NLP models in identifying framing bias, our study focuses on two primary categories of models: the GPT-4333GPT-4-0613 OpenAI (2023) and GPT-3.5-Turbo444GPT-3.5-turbo-0613 Ye et al. (2023), known for their state-of-the-art performance across a wide range of NLP benchmarks, and the open-source FLAN-T5 Wei et al. (2022), which, despite its accessibility, demonstrates significant efficacy in various tasks Chung et al. (2022). Additionally, we examined various sizes of the FLAN-T5 model, i.e., small (77M parameters), base (248M parameters), and large (783M parameters), to explore how model scale influences bias detection capabilities.
Our evaluation methodology employs a uniform series of prompts applied across all models under three distinct experimental conditions: (1) a zero-shot setting, probing the models’ inherent knowledge on framing bias detection; (2) a few-shot scenario, wherein models are provided with a small set of examples to inform their bias detection process; and (3) an explainability-focused approach, where models are asked to explain their rationale behind their decisions aside from label prediction.
Our evaluation framework is designed to thoroughly evaluate the models’ adeptness at navigating tasks without being finetuned on framing bias datasets. This aspect is particularly important within the domain of social science research, where specialised training datasets are often scarce.
3.2.1 Zero-Shot Prompting
In zero-shot settings, we deploy two distinct approaches to evaluate the models’ effectiveness in detecting framing bias. The first approach involves presenting the models with a straightforward task: determining if a headline is framed, without any additional context or examples.555This involves using the “Decide whether this claim is framed:” prompt for GPT models, and “Is this claim framed? OPTIONS Yes — No” for FLAN-T5. We chose the FLAN-T5 model’s prompt to ensure it aligns with the instructions that this model encountered during pretraining.
In the enhanced second zero-shot setting, we further add a specific definition to the input prompt to guide the task of framing detection. We use the “a communication strategy often used in journalism and political language, where certain aspects of an issue are highlighted while others are minimised or ignored, thereby promoting a particular interpretation of that issue” definition for this purpose.666This definition is grounded on the existing definitions of framing by Goffman (1974); Entman (1993, 2007). The model is then tasked with deciding if a given claim is framed based on this definition. This additional guidance aims to refine the models’ analytical lens, providing a clearer framework for identifying and evaluating framing within texts.
3.2.2 Few-Shot Prompting
In few-shot experiments, which build upon the zero-shot setting, we provide specific examples of both framed and not-framed headlines in the input prompt. These experiments are designed to investigate different factors influencing the model’s performance with example-based prompts:
Examining the Impact of Example Quantity:
To understand how the number of examples affects the model’s accuracy, we evaluate few-shot models in two different configurations. The first configuration includes a minimal set of two examples, one framed and one not framed. This setup helps us measure the baseline impact of including examples on model performance. The second configuration includes eight examples, from which four examples are framed headlines.
Assessing the Relevance of Examples:
We assess four scenarios to determine how the relevance of examples to the headline topics affects the model’s ability to detect framing:
Focused in-domain: In this setting, all the eight examples are relevant to gun violence, with a particular focus where all four framed headlines specifically address one aspect of gun violence news, which is health.
Varied in-Domain: In this setting framed examples cover four diverse aspects of gun violence777I.e., politics, public/school safety, race/ethnicity, and social/culture. to test the model’s adaptability to a range of in-domain cues.
Cross-Domain: To evaluate the model’s generalisation skills across topics, we use examples from completely different domains, such as immigration, in the few-shot prompts.
Mixed Domain: Combining in-domain and cross-domain examples, this scenario includes two framed instances related to gun violence and two from unrelated areas.
3.2.3 Explainable Prompting
To enhance our analysis, we revisited both the zero-shot and few-shot settings, introducing an additional requirement: the model must not only determine if a claim is framed but also provide an explanation for its decision. To achieve this, we appended the instruction “then give an explanation for your response” to the prompts.
| GPT-3.5 Turbo | GPT-4 | ||||||||
| Explainable | Explainable | ||||||||
| F | Acc. | F | Acc. | F | Acc. | F | Acc. | ||
| Zero-Shot | No Definition | 20.48 | 52.70 | 61.43 | 53.28 | 64.96 | 58.40 | 63.75 | 60.41 |
| +Definition | 51.55 | 59.14 | 59.03 | 58.79 | 65.36 | 63.84 | 64.64 | 60.99 | |
| Few-Shot | 2 examples | 26.83 | 54.16 | 65.84 | 61.68 | 58.25 | 65.84 | 64.23 | 64.92 |
| 8 (focused in-domain) | 40.72 | 58.25 | 65.32 | 61.57 | 64.50 | 63.30 | 66.44 | 63.96 | |
| 8 (varied in-domain) | 46.36 | 60.22 | 64.40 | 60.79 | 68.51 | 65.38 | 70.41 | 66.92 | |
| 8 (cross domain) | 41.22 | 57.67 | 60.76 | 62.08 | 59.42 | 66.04 | 59.09 | 64.61 | |
| 8 (mixed domain) | 56.93 | 63.49 | 64.60 | 62.99 | 63.33 | 64.37 | 63.87 | 63.76 | |
| FLAN-T5 Small | FLAN-T5 Base | FLAN-T5 Large | |||||||||||
| Explainable | Explainable | Explainable | |||||||||||
| F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | ||
| Zero-Shot | No Definition | 0.46 | 49.06 | 17.96 | 49.21 | 0 | 0 | 25.86 | 52.74 | 34.33 | 46.19 | 51.01 | 49.06 |
| +Definition | 0.15 | 50.13 | 1.21 | 50.10 | 0 | 0 | 0.61 | 50.06 | 42.27 | 48.06 | 37.63 | 43.96 | |
| Few-Shot | 2 examples | 0.31 | 50.06 | 0.31 | 50.06 | 0 | 0 | 3.12 | 49.98 | 42.27 | 48.06 | 37.63 | 43.96 |
| 8 (focused in-domain) | 0.15 | 50.17 | 3.84 | 50.10 | 0 | 0 | 4.83 | 50.17 | 16.78 | 47.53 | 36.85 | 44.69 | |
| 8 (varied in-domain) | 0 | 0 | 0.15 | 49.94 | 0 | 0 | 1.66 | 49.94 | 22.78 | 45.96 | 35.03 | 45.27 | |
| 8 (cross domain) | 5.30 | 49.37 | 2.95 | 49.60 | 0 | 0 | 4.99 | 50.33 | 40.23 | 45.11 | 51.64 | 46.88 | |
| 8 (mixed domain) | 0.31 | 50.21 | 1.51 | 50.02 | 0 | 0 | 3.84 | 50.13 | 34.39 | 44.58 | 44.05 | 43.43 | |
4 LLMs’ Reliability in Detecting Framing Bias
The evaluation results for GPT and FLAN-T5 models across different settings are presented in Table 1 and Table 2, respectively. We can summarise the notable findings of the results of as follows:
Explainable Prompting Enhances Reliability:
Explainable prompting consistently yielded more reliable outcomes in both zero- and few-shot variations, as shown by the reduced variance in accuracy and F scores. This underscores the importance of explainable settings for social science research on framing bias.
Optimal Performance in Few-Shot with Diverse Examples:
The GPT-4 model achieved the highest accuracy and F scores in few-shot scenarios with a wide range of in-domain examples (8 varied in-domain). This finding suggests that incorporating diverse, relevant in-domain examples significantly improves model performance when the news headline domain is known.
Tendency Towards Framing Bias Classification:
In explainable settings, both GPT-3.5 and GPT-4 models exhibit a higher tendency to classify headlines as framed. This is indicated by higher F scores for framed headline detection alongside lower overall accuracy, suggesting a bias towards classifying headlines as framed when prompted to identify framing.
Challenges with Cross-Domain Examples:
F scores dropped in cross-domain settings without explainable prompts, falling below those in zero-shot settings with definitions. This highlights the challenges of applying advanced models to new or mixed domains, where headline topics vary widely. Therefore, for new domain applications, we suggest to either employ a zero-shot setting with task definition, or opting for explainable prompts when examples are used in the input prompts. However, for large-scale data analysis, opting for a zero-shot approach without explanations could be more efficient, albeit at the potential cost of reduced reliability compared to the explainable setting.
Inherent Challenge in Framing Bias Detection:
Across both models, the highest F scores observed were in the range of 60-70 points, highlighting the inherent difficulty of detecting framing bias. Framing detection requires understanding subtle language nuances, contextual cues, and the intended message framing, which are complex and context-dependent tasks. Furthermore, the performance between GPT-3.5 and GPT-4 did not significantly differ, particularly in more realistic use cases like zero-shot and cross-domain settings. Considering the higher cost of using GPT-4’s API, GPT-3.5 might be a more cost-effective choice for framing bias detection in large scale analysis.
FLAN-T5 Performance Lag:
The results from Table 2 show that FLAN-T5 models significantly underperform compared to GPT models, with the best F score reaching only 51 points. This indicates that smaller models like FLAN-T5 may not yet be viable for complex tasks such as framing bias detection without additional task-specific fine-tuning or more extensive training data.
5 Analysis
In this section, we perform a set of additional analysis and experiments to identify potential biases and areas for future research enhancements.
| Headline: This list of lives lost in mass shootings grows longer. Here are their names. |
| Gold: Not Framed, GPT-4: Framed, GPT-3.5: Not Framed |
| GPT-4 Explanation: This claim is framed because it specifically focuses on the tragic aspect of lives lost in mass shootings, evoking empathy and sadness. |
| While it’s factual that people have died, the phrasing and focus frame the issue in a particular emotional light. |
| GPT-3.5 Explanation: This claim is not framed. It simply states that the list of lives lost in mass shootings is growing longer and presents the names of the victims. |
| It does not provide any particular perspective or agenda. |
| Headline: Pregnant woman who killed intruder in justified shooting now faces felony gun charge due to previous marijuana conviction. |
| Gold: Not Framed, GPT-4: Framed, GPT-3.5: Not Framed |
| GPT-4 Explanation: The claim is framed. It selectively emphasizes on the individual’s status as a pregnant woman and the nature of the shooting |
| as justified, while sidelining the legal issue of her previous marijuana conviction which is directly linked to the current felony charge. |
| The framing encourages empathy for the person involved by prominently highlighting her pregnancy and the self-defense aspect, provoking an emotional response. |
| GPT-3.5 Explanation: This claim appears to accurately represent the situation. It states that a pregnant woman killed an intruder in a justified |
| shooting and is now facing a felony gun charge due to a previous marijuana conviction. The claim does not seem to contain any misleading or false information. |
| Headline: Gunman upset by divorce is found dead after manhunt. |
| Gold: Not Framed, GPT-4: Framed, GPT-3.5: Not Framed |
| GPT-4 Explanation: This claim can be considered as framed because it specifically points out the gunman’s divorce as a possible motive. |
| By doing so, the author is driving the audience’s view to place emphasis on the gunman’s personal issues, making them potentially more sympathetic to him. |
| Other potential factors or the bigger context are not discussed here. |
| GPT-3.5 Explanation: This claim appears to be accurately presented. |
| It provides a straightforward account of events, stating that a gunman who was upset by a divorce was found dead after a manhunt. |
| There is no apparent bias or manipulation in the framing of this claim. |
The Impact of Emotional Language on Framing Bias Detection
Our analysis of GPT-4’s errors reveals a consistent pattern: the model frequently interprets emotional language as an indicator of framing bias, irrespective of the headline’s actual framing. This tendency likely stems from the observation that emotionally charged language often accompanies framed headlines, evoking specific frames and eliciting emotional responses from readers, thereby influencing their perception of the news narrative Valkenburg et al. (1999). However, this leads to a biased error in the model, mistakenly identifying expressions of emotion as indicative of framing bias. Such a pattern highlights the challenge of distinguishing between genuine emotional expression and intentional framing bias. Making this distinction is essential for enhancing the model’s precision in analysing news narratives. Table 3 provides examples of such misclassifications by GPT-4 in the explainable zero-shot setting. On the other hand, GPT-3.5 seems to be less affected by the emotional language.
| GPT-3.5 | GPT-4 | Flan-T5 | ||||||||
| Small | Base | Large | ||||||||
| F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | |
| Clear | 71.43 | 70.15 | 75.81 | 77.61 | 3.23 | 42.86 | 9.23 | 43.81 | 49.6 | 40.00 |
| Contested | 41.43 | 38.81 | 3.60 | 20.15 | 3.03 | 39.62 | 0 | 0 | 65.60 | 59.43 |
| Clear | Contested | |
| Agreement | 69.40 | 58.95 |
| Agreement GL | 58.20 | 16.45 |
| GPT-3.5 | GPT-4 | |||||||
| Explainable | Explainable | |||||||
| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | |
| No Definition | 20.48 | 52.70 | 61.43 | 53.28 | 64.96 | 58.40 | 63.75 | 60.41 |
| Different Wording | 25.39 | 53.55 | 59.04 | 54.64 | 47.40 | 60.56 | 45.55 | 57.76 |
| +Definition | 51.55 | 59.14 | 59.03 | 58.79 | 65.36 | 63.84 | 64.64 | 60.99 |
| Different Wording | 58.64 | 58.16 | 63.56 | 59.44 | 57.03 | 62.63 | 59.02 | 61.07 |
| Framed | Not-Framed | |
| GPT-3.5 Turbo | 84 | 38 |
| GPT-4 | 108 | 22 |
| Flan-T5 Small | 1 | 129 |
| Flan-T5 Base | 0 | 130 |
| Flan-T5 Large | 32 | 98 |
Clear vs. Contested Cases of Framing Annotations
During our initial analysis, we observed discrepancies in some data annotations, with some annotations being potentially incorrect. Additionally, there were ambiguous cases where multiple labels could be justified based on different interpretations. In this regard, we manually examined half of the headlines in GVFC, i.e., 1300 headlines. Within this subset, we identified 134 contested annotations. Examples include “Live: Trump visits Pittsburgh after synagogue shooting” and “Shopify bans sale of certain firearms, accessories”, both annotated as framed. However, not all 134 cases are necessarily wrong annotations. Some of them, such as “Thousands gather to honor victims of the mass shooting with tears, candlelight, and song”, present more nuanced challenges in framing detection.88863.4% of headlines in this subset are annotated as framed in the dataset.
Additionally, we selected another set of 134 headlines where the presence or absence of framing bias was more apparent. For instance, headlines such as “Two dead including shooter at Florida yoga studio” and “Parkland school shooter blames massacre on a ’demon’ voice” serve as clear examples of non-framed and framed headlines that are also annotated correctly, respectively. We call this subset the “Clear” subset.
Table 4 reports the results of evaluating our models on these two distinct subsets. As expected, both GPT models exhibited strong performance on the Clear subset. However, they have a considerably low results on the contested subset, with GPT-4’s F score dropping to 3.6 points.
We have also calculated how often GPT-4 and GPT-3.5’s predictions aligned within these subsets. These agreement ratios are shown in Table 5. We observe that while the models reached an agreement of about 59% on the contested subset, only 16% of these agreed-upon predictions matched the gold labels. This suggests that these models’ can be useful in flagging potential annotation inaccuracies within existing or new datasets. As an example, both models identify the “ Muslim Americans raise more than $200,000 for those affected by Pittsburgh synagogue shooting” as not framed while it is annotated as framed in the data.
Impact of Prompt Variability
One of the challenges with prompt-based techniques is their sensitivity to the wording of input prompts; even minor variations can lead to drastically different performances. We examined this variability by using various input examples in focused in-domain, varied in-domain, cross-domain, and mixed-domain settings as shown in Table 1. To further examine the impact of prompt wording on framing bias detection results, we altered the prompt from “Decide whether this claim is framed or not” to “Decide whether this claim contains framing bias” in our zero-shot experiments. Table 6 shows these findings. Our result indicates that the explainable setting yields the most consistent results, where different formulations of the question minimally affects outcome variability. On the other hand, we observe the most significant performance change in the GPT-4 model without a defined task or explanation, where the F score drops from 64 to 47 points. Given that the explainable setting requires more computational resources, including a task definition is a more efficient alternative for decreasing variability in results, and therefore enhancing their reliability.
Evaluation in the Wild
Until now, our results and analysis were centered on the GVFC dataset, which is limited to the narrow subject of gun violence within the U.S. To assess the models in a scenario closer to real-world conditions, where the topics of headlines are varied and unknown, we collected a new dataset from a website known for its emphasis on news framing.999https://newsframes.wordpress.com/category/headlines/ This collection consists of 130 headlines spanning diverse subjects such as British weather, health, and European matters.
We assessed the models on this novel dataset employing the explainable few-shot setting with 8 cross-domain examples, which resulted in reliable results in Section 4. Table 7 shows the counts of headlines predicted as framed or not framed by each model. Given the dataset’s origin from a website known for featuring framed headlines, we anticipated the majority of headlines to be framed. We observe that the GPT-4 model predominantly identified headlines as framed. However, considering GPT-4’s tendency towards classifying inputs as framed, a promising avenue for future research involves developing a dataset specifically aimed at challenging the models with not-framed cases. This could include not-framed headlines that still incorporate emotional language, as discussed in the beginning of this section.
6 Conclusions
This work advances our understanding of the performance of LLMs such as GPT-3.5 Turbo, GPT-4, and Flan-T5 (small, base, and large) in detecting framing bias within news headlines. Through a comprehensive experimental approach that included zero-shot, few-shot, and explainable prompting settings, we found that GPT-4 excels in few-shot scenarios with a variety of in-domain examples, highlighting its superior capability in recognising framing bias. However, GPT-4 also displayed a tendency to classify headlines as framed, suggesting a valuable line of research could involve testing these models on datasets containing non-framed headlines that potentially also incorporate emotional language, identified as another potential weakness. In fact, our research explores the impact of emotional language on framing bias detection, with GPT-3.5 showing resilience against emotional language.
The study also identifies a performance shortfall in Flan-T5, underscoring the challenges smaller models face in complex detection tasks without specific fine-tuning. Our findings indicate that F1 scores drop in cross-domain scenarios with non-explainable prompts, suggesting the effectiveness of zero-shot approaches with clear definitions or explainable prompts for new domain applications. The study further demonstrates the utility of these models in identifying potential annotation inaccuracies in new or already existing datasets.
7 Limitations
The findings of this study have to be seen in light of some limitations. For example, our evaluation focuses solely on English-language content, leaving space for further investigation on other languages to explore our findings’ applicability to non-English contexts. This limitation suggests a need for further investigation into the performance of LLMs across different languages and cultural contexts to fully assess the potential use of these models in social science research for detecting framing bias and analysing media narratives.
Furthermore, two of the five models evaluated in our work are accessible only through OpenAI’s API, which is closed-source and subject to changes over time. This could affect the reproducibility of our results with newer versions of API, and they may have their own limitations. Therefore, focusing on improving open-source models emerges as a critical pathway forward, ensuring broader accessibility and reproducibility in research.
8 Acknowledgements
This work was supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1].

### v3

@@
-Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Bias Detection
+Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Detection
 Abstract
-This work contributes to the expanding research on the applicability of LLMs in social sciences by examining the performance of GPT-3.5 Turbo, GPT-4, and Flan-T5 models in detecting framing bias in news headlines through zero-shot, few-shot, and explainable prompting methods.
-A key insight from our evaluation is the notable efficacy of explainable prompting in enhancing the reliability of these models, highlighting the importance of explainable settings for social science research on framing bias.
-GPT-4, in particular, demonstrated enhanced performance in few-shot scenarios when presented with a range of relevant, in-domain examples.
-FLAN-T5’s poor performance indicates that smaller models may require additional task-specific fine-tuning for identifying framing bias detection.
-Our study also found that models, particularly GPT-4, often misinterpret emotional language as an indicator of framing bias, underscoring the challenge of distinguishing between reporting genuine emotional expression and intentionally use framing bias in news headlines.
-We further evaluated the models on two subsets of headlines where the presence or absence of framing bias was either clear-cut or more contested, with the results suggesting that these models’ can be useful in flagging potential annotation inaccuracies within existing or new datasets.
-Finally, the study evaluates the models in real-world conditions ("in the wild"), moving beyond the initial dataset focused on U.S.
-Gun Violence, assessing the models’ performance on framed headlines covering a broad range of topics111All data created in this study will be made publicly available..
-Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Bias Detection
+Previous studies on framing have relied on manual analysis or fine-tuning models with limited annotated datasets.
+However, pre-trained models, with their diverse training backgrounds, offer a promising alternative.
+This paper presents a comprehensive analysis of GPT-4, GPT-3.5 Turbo, and FLAN-T5 models in detecting framing in news headlines.
+We evaluated these models in various scenarios: zero-shot, few-shot with in-domain examples, cross-domain examples, and settings where models explain their predictions.
+Our results show that explainable predictions lead to more reliable outcomes.
+GPT-4 performed exceptionally well in few-shot settings but often misinterpreted emotional language as framing, highlighting a significant challenge.
+Additionally, the results suggest that consistent predictions across multiple models could help identify potential annotation inaccuracies in datasets.
+Finally, we propose a new small dataset for real-world evaluation on headlines from a diverse set of topics.111All data created in this study will be made publicly available.
+Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Detection
 Valeria Pastorino, Jasivan A. Sivakumar, Nafise Sadat Moosavi Department of Computer Science University of Sheffield United Kingdom {vpastorino1|jasivakumar1|n.s.moosavi}@sheffield.ac.uk
 1 Introduction
-In today’s digital age, the proliferation of news sources and the rapid dissemination of information highlight the essential need for unbiased reporting.
-Framing bias, which manipulates news content to potentially shift public perception, presents a formidable obstacle in maintaining a well-informed and unbiased public sphere Binotto and Bruno (2018).
-This form of bias not only influences how events are portrayed but also impacts public attitudes and policy decisions.
-For instance, the reporting on a government policy change might be emotionally charged in one headline as “Government’s heartless cutbacks leave thousands without essential services”, whereas another might offer a more neutral perspective: “Government announces reduction in funding for public services”.
-The concept of framing has received extensive scrutiny in the social sciences, identified as a technique that selectively highlights certain realities to shape audience perceptions and reactions Goffman (1974); Entman (1993).
-While its significant effect on public communication and media is well-documented, the analysis of framing has historically been challenged by its fragmented nature and reliance on manual, small-scale studies.
-This situation underscores the necessity for new methods capable of dissecting the complexities of framing in media comprehensively.
-Amid this landscape, with an overwhelming array of news sources and headlines, it becomes crucial to determine whether they convey biased messages, intentionally or inadvertently.
-This determination is not merely academic but essential for preserving the integrity of public discourse.
-The introduction of large language models (LLMs) opens up novel pathways for identifying framing bias, especially in news headlines.
-The accessibility and advanced capabilities of models, particularly those within the GPT series, have garnered attention from professionals across various fields, including media analysis and political science, seeking to navigate the extensive corpus of available information.
-However, despite their impressive success across many applications, the reliability of these models in executing nuanced tasks like detecting framing bias remains an open question.
-In this study, we address this gap by providing an in-depth investigation into the strengths and shortcomings of state-of-the-art natural language processing (NLP) models in identifying framing bias.
-We perform a comprehensive examination of the performance of two top-tier NLP models, GPT-4 and GPT-3.5, alongside the open-source FLAN-T5 model (Wei et al., 2022), which has a reasonable performance across different NLP tasks.
-Our findings reveal interesting patterns: while the pre-trained FLAN-T5 struggles with this task, the GPT models show promising results.
-Our results suggest that prompting the models to explain their decisions results in more reliable performances across different settings.
-A standout discovery is the enhanced reliability of predictions when models are prompted to explain their reasoning, a technique that also narrows the variability in outcomes across different example sets.
-To test the real-world applicability of these models, apart from a standard dataset that is annotated by experts, we automatically collect a new evaluation set containing framed headlines from different domains.
-The GPT models also accurately identifying the majority of framed titles in this dataset.
-Our analysis suggests that while the examined GPT models achieve high scores on the examined evaluation sets, there are still areas that they struggle.
-As an example, GPT-4 often mistook emotional language for framing bias.
-This insight points to the need for developing more nuanced test sets that can present more complex challenges for evaluation.
-Additionally, our experiments suggest a potential use case for these models beyond mere bias detection.
-When GPT models, particularly in a more reliable prompting setting, converge on a prediction that deviates from the established gold standard, it may hint at underlying annotation inaccuracies within the dataset.
-Our findings suggest that a future direction involves constructing more challenging datasets that mirror the real-world complexities of framing bias, as well as enhancing the performance of smaller and open-source models on this challenging task.
+In today’s digital age, the spread of news sources and the rapid dissemination of information highlight the essential need for unbiased reporting.
+Framing is a communication strategy used to manipulate news content in order to potentially shift public perception to follow a certain agenda, and is an obstacle in maintaining a well-informed and unbiased public audience Binotto and Bruno (2018).
+Framing not only influences how events are perceived and remembered but can also impact overall attitudes and policy decisions, since it helps building a potentially biased narrative around an event.
+For instance, the reporting on a government policy change might be framed in one headline as “Government’s heartless cutbacks leave thousands without essential services”, whereas another might offer a more neutral perspective: “Government announces reduction in funding for public services”.
+The concept of framing has received extensive analysis in social science, identified as a technique that selectively highlights certain realities to shape audience perceptions and reactions Goffman (1974); Entman (1993).
+Despite its well-documented impact on public communication and media, analysing framing has traditionally been limited by manual, small-scale studies Baumer et al.
+(2015).
+This limitation highlights the urgent need for new methods capable of thoroughly analysing the complexities of framing in media.
+In today’s context, with an overwhelming array of news sources and headlines, it is crucial to determine whether they convey framed messages.
+This task is not merely academic but essential for preserving the integrity of public discourse.
+The introduction of large language models (LLMs) provides a promising alternative for identifying framing bias in news headlines.
+LLMs, already trained on vast corpus of text data, can be readily adapted or tested for similar tasks using prompts or a few examples.
+Despite their success in many areas, their reliability in handling nuanced tasks like detecting framing remains an open question.
+This study aims to address this gap by providing an in-depth investigation into the strengths and shortcomings of prominent instruction-following pre-trained language models in identifying framing in news headlines.
+We perform a comprehensive examination of the performances of GPT-4, GPT-3.5 Turbo, and open source FLAN-T5 model.
+We chose FLAN-T5 due to its robust performance across various NLP tasks (Wei et al., 2022) and its status as one of the first models designed to follow instructions or prompts effectively.
+Our results reveal that, while out-of-the-box FLAN-T5 models struggle with framing detection, the GPT models show promising results.
+Notably, prompting the models to explain their decisions results in more reliable performance across different settings.
+However, our analysis indicates that while the GPT models perform well on the evaluation sets, they still face some challenges.
+For instance, GPT-4 often mistook emotional language for framing.
+This insight underscores the need for more nuanced framing test sets that present complex challenges for evaluation.
+Additionally, our experiments suggest a potential use case for these models beyond mere framing detection.
+In fact, by dividing the dataset into clear-cut and contested subsets based on the original annotations, we observed that when GPT models, particularly in a more reliable prompting setting, agree in deviating from the established gold standard, it may indicate underlying annotation inaccuracies.
+To assess the real-world applicability of these models across various domains, we collected a new evaluation set containing framed headlines from different topics, achieving an accuracy of 75.16% for GPT-3.5 Turbo and 80.25% for GPT-4.
+Finally, our study shows that, when fine-tuned, FLAN-T5 outperforms both GPT models on in-domain data (GVFC) with an accuracy of 78.83% compared to GPT-4’s 66.92%.
+However, this advantage diminishes with out-of-domain evaluation in real-world settings, where its accuracy drops to 68.78% compared to GPT-4’s 80.25%.
+This emphasises the effectiveness of prompt-based techniques without fine-tuning for analysing the presence of framing in real world scenarios.
 2 Related Work
 2.1 Automatic Framing Detection
-Various NLP methodologies have been used to automate framing bias detection.
-Early studies in framing detection primarily utilized Topic Modeling, Structural Topic Modeling, and Hierarchical Topic Modeling to uncover themes and topics within large datasets DiMaggio et al.
+Various NLP methodologies have been used to automate framing detection.
+Early studies in framing detection primarily utilised Topic Modeling, Structural Topic Modeling, and Hierarchical Topic Modeling to uncover themes and topics within large datasets DiMaggio et al.
 (2013); Nguyen et al.
@@
 (2021) for computational framing analysis.
-However, as noted by Ali and Hassan (2022), the emphasis in such approaches remains on detecting topics rather than the nuanced framing of those topics.
-The focus on topics also derives from the connections between agenda setting and framing strategies in computational social sciences, with studies analysing these two phenomena together Field et al.
+However, as noted by Ali and Hassan (2022), the emphasis in such approaches remains on detecting topics rather than the nuanced framing of those events.
+The focus on topics could also derive from the connections between agenda setting and framing strategies in computational social sciences, with studies analysing these two phenomena together Field et al.
 (2018).
@@
 Tourni et al.
-(2021) demonstrated that combining transformers models for processing news headlines with residual network models to process news lead images could lead to accurate framing bias detection.
-Similarly, Naderi and Hirst (2017) explored the use of various deep neural networks, such as LSTMs, BiLSTMs, and GRUs, for frame prediction at the sentence level using the Media Frame Corpus (MFC) Card et al.
+(2021) demonstrated that combining transformer models for processing news headlines with residual network models to process news lead images could improve the accuracy of framing detection.
+Similarly, Naderi and Hirst (2017) explored the use of various neural networks, such as LSTMs, BiLSTMs, and GRUs, for frame prediction at the sentence level using the Media Frame Corpus (MFC) Card et al.
 (2015).
@@
 (2019) to predict frames in news headlines.
-Their work resulted in the creation of the Gun Violence Frame Corpus (GVFC), a benchmark dataset for framing analysis which will be further discussed in section 2.1.1.
+Their work resulted in the creation of the Gun Violence Frame Corpus (GVFC), a benchmark dataset for framing analysis which will be further discussed in section 2.2.
 More recently, in a novel approach to automatic framing detection, Lai et al.
 (2022) introduced an unsupervised learning method leveraging Wikipedia’s category system to identify frames within news articles, using the GVFC to evaluate this method.
-2.1.1 Datasets
+2.2 Datasets
 Media Frame Corpus (MFC)
 MFC (Card et al., 2015) is a collection of annotated U.S.
-newspaper articles on topics like immigration, smoking, and same-sex marriage, analyzed for framing.
+newspaper articles on topics like immigration, smoking, and same-sex marriage, analysed for framing.
 Utilising the Policy Frames Codebook (PFC) by Boydstun et al.
@@
 Despite achieving an inter-coder reliability (ICR) of 0.60, critiques, particularly from Ali and Hassan (2022), argue that the PFC’s broad dimensions conflate topics with frames, potentially missing nuanced strategic framing.
-Moreover, MFC categorises content into wide-ranging dimensions that might not always precisely capture the specific framing intended by a news headline.
+Moreover, MFC categorises content into wide-ranging dimensions (i.e., politics, economic, etc.) that might not always precisely capture the specific framing intended by a news headline Ali and Hassan (2022).
 This categorisation can make it difficult to directly identify whether and how a headline is framed without a deeper, nuanced analysis.
+For example, the headline “Where the jobs are” has been annotated using the framing dimension “Economic”, while the following portion of text “Critics of illegal immigration can make many cogent arguments to support the position that the U.S.
+Congress and the Colorado legislature must develop effective and well-enforced immigration policies that will restrict the number of people who migrate here legally and illegally.” was annotated using the framing dimension “Policy prescription”.
 Gun Violence Frame Corpus (GVFC)
-Another significant dataset in the field of framing analysis is the Gun Violence Frame Corpus (GVFC), introduced by Liu et al.
+Another significant dataset in the field of framing analysis is the GVFC, introduced by Liu et al.
 (2019).
@@
 A specialised codebook was then developed, serving as a training tool for annotators along with annotation guidelines.
-GVFC is made of 2990 news headlines, with 1,300 headlines specific to the issue of Gun Violence in the United States.
-All the headlines are coded to have a primary frame, while only 319 have 2 frames.
+GVFC is made of 2990 news headlines, with 2616 headlines specific to the issue of Gun Violence in the United States.
+All the in-domain headlines are coded to have a primary frame, while only 319 have two frames.
 For instance, the headline “It’s Time to Hand the Mic to Gun Owners” is annotated with “Public opinion” as the first frame and “2nd Amendment” as the second frame.
-Similarly, “Trevor Noah: The Second Amendment Is Not Intended for Black People” is annotated with “2nd Amendment” and “Race/Ethnicity” frames (Liu et al., 2019, p.
-507).
+Similarly, “Trevor Noah: The Second Amendment Is Not Intended for Black People” is annotated with “2nd Amendment” and “Race/Ethnicity” frames (Liu et al., 2019).
 Non-English Data
@@
 gun violence.
-This extension involved training two native speakers per language to annotate headlines—350 in German, 200 in Turkish, and 210 in Arabic.
+This extension involved training two native speakers per language to annotate headlines - 350 in German, 200 in Turkish, and 210 in Arabic.
 Piskorski et al.
-(2023b) presented an annotated dataset made of articles spanning though nine languages: English, French, German, Georgian, Greek, Italian, Polish, Russian, and Spanish.
+(2023b) presented an annotated dataset made of articles spanning through nine languages: English, French, German, Georgian, Greek, Italian, Polish, Russian, and Spanish.
 This dataset addresses a variety of topics including the COVID-19 pandemic, abortion-related legislation, migration, Russo-Ukrainian war, and various parliamentary elections.
@@
 (2023a).
-While our study focuses on the English language, exploring framing bias detection in less resourced languages presents a promising future direction.
-By establishing a foundation in English, we lay the groundwork for future research to investigate the strengths and shortcomings of framing bias detection in languages beyond English.
 2.3 Evaluating LLMs in Social Science
@@
 (2023), and controversial issues and bias Sun et al.
-(2023), has received increasing interest, showcasing a wide range of strengths and limitations unique to each task.
+(2023), has received increasing interest, showcasing a wide range of strengths and limitations of the examined language models unique to each task.
 This diversity stems from the specific challenges and nuances of social phenomena.
 Although LLMs excel in generating and understanding human-like text, the complex requirements of social science tasks necessitate a detailed, task-specific examination of their performance and reliability.
-In this work, we contribute to the expanding research on the applicability of LLMs in social sciences by specifically investigating their reliability in detecting framing bias.
-This exploration not only aims to identify the strengths and shortcomings of state-of-the-art models in a crucial area of framing analysis but also to extend our comprehension of how these models can be optimised for the nuanced task of framing detection.
+In this work, we contribute to the expanding research on the applicability of LLMs in social sciences by specifically investigating their reliability in detecting framing.
 3 Experimental Setup
 3.1 Data
-For our evaluation, we selected the Gun Violence Framing Corpus (GVFC) dataset Liu et al.
+For our evaluation, we selected GVFC Liu et al.
 (2019), motivated by its comprehensive coverage of U.S.
-Gun Violence framing through 2900 annotated newspaper headlines as well as the high ICR met in the annotation process.222In our experiments, we have excluded the headlines that are not relevant to Gun Violence.
+Gun Violence framing as well as the high ICR met in the annotation process.
+In our experiments, we have excluded the headlines that are not relevant to Gun Violence and hence, are not annotated with framing information.
+The dataset’s annotations identify whether each headline reflects any of nine critical aspects of gun violence framing: gun rights, gun control, politics, mental health, public/school safety, race/ethnicity, public opinion, social/cultural issues, and economic consequences.
+We consider headlines tagged with any of the above categories as framed, while all others as not framed.
 Furthermore, we have excluded 22 relevant headlines in order to use them for Few-Shot in domain prompting, leaving 2594 relevant headlines for our analysis.
-The dataset’s annotations identify whether each headline reflects any of nine critical aspects of gun violence framing: gun rights, gun control, politics, mental health, public/school safety, race/ethnicity, public opinion, social/cultural issues, and economic consequences.
-In our study, headlines tagged with any of these framing aspects were categorised as framed, while all others were classified as not framed.
+Of these, 1293 are framed and 1301 are not framed.
+Hence, the majority label is Not Framed and the majority baseline is 50.15%.
 3.2 Models
-To assess the capabilities of state-of-the-art NLP models in identifying framing bias, our study focuses on two primary categories of models: the GPT-4333GPT-4-0613 OpenAI (2023) and GPT-3.5-Turbo444GPT-3.5-turbo-0613 Ye et al.
-(2023), known for their state-of-the-art performance across a wide range of NLP benchmarks, and the open-source FLAN-T5 Wei et al.
+To assess the capabilities of two of the leading closed-source models in identifying framing, our study focuses on the GPT-4222GPT-4-0613.
+Temperature parameter set to 0.
+OpenAI (2023) and GPT-3.5-Turbo333GPT-3.5-turbo-0613.
+Temperature parameter set to 0.
+Ye et al.
+(2023), known for their state-of-the-art performance across a wide range of NLP benchmarks.
+We chose GPT-4 and GPT-3.5 Turbo due to their advanced capabilities and ease of use.
+While they are closed-source with a paid API, they are easily accessible for social scientists, eliminating the need for setting up libraries and running models with high memory and expertise requirements.
+This accessibility makes them practical tools for researchers who may not have extensive computational resources.
+Additionally, we chose the open-source FLAN-T5 Wei et al.
 (2022), which, despite its accessibility, demonstrates significant efficacy in various tasks Chung et al.
 (2022).
-Additionally, we examined various sizes of the FLAN-T5 model, i.e., small (77M parameters), base (248M parameters), and large (783M parameters), to explore how model scale influences bias detection capabilities.
-Our evaluation methodology employs a uniform series of prompts applied across all models under three distinct experimental conditions: (1) a zero-shot setting, probing the models’ inherent knowledge on framing bias detection; (2) a few-shot scenario, wherein models are provided with a small set of examples to inform their bias detection process; and (3) an explainability-focused approach, where models are asked to explain their rationale behind their decisions aside from label prediction.
-Our evaluation framework is designed to thoroughly evaluate the models’ adeptness at navigating tasks without being finetuned on framing bias datasets.
+We examined various sizes of the FLAN-T5 model, i.e., small (77M parameters), base (248M parameters), and large (783M parameters), to explore how model scale influences framing detection capabilities.
+Our evaluation methodology employs a uniform series of prompts applied across all models under three distinct experimental conditions: (1) a zero-shot setting, probing the models’ inherent knowledge on framing detection; (2) a few-shot scenario, wherein models are provided with a small set of examples to inform their framing detection process; and (3) an explainability-focused approach, where models are asked to explain their rationale behind their decisions aside from label prediction.
+Our evaluation framework is designed to thoroughly evaluate the models’ adeptness at navigating tasks without being fine-tuned on framing datasets.
 This aspect is particularly important within the domain of social science research, where specialised training datasets are often scarce.
-3.2.1 Zero-Shot Prompting
-In zero-shot settings, we deploy two distinct approaches to evaluate the models’ effectiveness in detecting framing bias.
-The first approach involves presenting the models with a straightforward task: determining if a headline is framed, without any additional context or examples.555This involves using the “Decide whether this claim is framed:” prompt for GPT models, and “Is this claim framed?
-OPTIONS Yes — No” for FLAN-T5.
-We chose the FLAN-T5 model’s prompt to ensure it aligns with the instructions that this model encountered during pretraining.
+3.3 Zero-Shot Prompting
+In zero-shot settings, we deploy two distinct approaches to evaluate the models’ effectiveness in detecting framing.
+The first approach involves presenting the models with a straightforward task: determining if a headline is framed, without any additional context or examples.444GPT prompt: “Decide whether this claim is framed”.
+FLAN-T5 prompt:“Is this claim framed? OPTIONS Yes — No”. We chose the FLAN-T5 model’s prompt to ensure it aligns with the instructions that this model encountered during pre-training.
 In the enhanced second zero-shot setting, we further add a specific definition to the input prompt to guide the task of framing detection.
-We use the “a communication strategy often used in journalism and political language, where certain aspects of an issue are highlighted while others are minimised or ignored, thereby promoting a particular interpretation of that issue” definition for this purpose.666This definition is grounded on the existing definitions of framing by Goffman (1974); Entman (1993, 2007).
+We use the “a communication strategy often used in journalism and political language, where certain aspects of an issue are highlighted while others are minimised or ignored, thereby promoting a particular interpretation of that issue” definition for this purpose.555This definition is grounded on the existing definitions of framing by Goffman (1974); Entman (1993, 2007).
 The model is then tasked with deciding if a given claim is framed based on this definition.
 This additional guidance aims to refine the models’ analytical lens, providing a clearer framework for identifying and evaluating framing within texts.
-3.2.2 Few-Shot Prompting
+3.4 Few-Shot Prompting
 In few-shot experiments, which build upon the zero-shot setting, we provide specific examples of both framed and not-framed headlines in the input prompt.
@@
 To understand how the number of examples affects the model’s accuracy, we evaluate few-shot models in two different configurations.
-The first configuration includes a minimal set of two examples, one framed and one not framed.
+The first configuration includes a minimal set of two GVFC examples, one framed and one not framed.
 This setup helps us measure the baseline impact of including examples on model performance.
@@
 Assessing the Relevance of Examples:
-We assess four scenarios to determine how the relevance of examples to the headline topics affects the model’s ability to detect framing:
+We assess four scenarios to determine how the relevance of the examples to the headline topics in the test set, if known a priori, affects the model’s ability to detect framing:
 Focused in-domain: In this setting, all the eight examples are relevant to gun violence, with a particular focus where all four framed headlines specifically address one aspect of gun violence news, which is health.
-Varied in-Domain: In this setting framed examples cover four diverse aspects of gun violence777I.e., politics, public/school safety, race/ethnicity, and social/culture.
+Varied in-Domain: In this setting framed examples cover four diverse aspects of gun violence666I.e., politics, public/school safety, race/ethnicity, and social/culture.
 to test the model’s adaptability to a range of in-domain cues.
-Cross-Domain: To evaluate the model’s generalisation skills across topics, we use examples from completely different domains, such as immigration, in the few-shot prompts.
+Cross-Domain: To evaluate the model’s performance on topics not known in advance, we use examples from completely different domains, such as immigration.
 Mixed Domain: Combining in-domain and cross-domain examples, this scenario includes two framed instances related to gun violence and two from unrelated areas.
-3.2.3 Explainable Prompting
+3.5 Explainable Prompting
 To enhance our analysis, we revisited both the zero-shot and few-shot settings, introducing an additional requirement: the model must not only determine if a claim is framed but also provide an explanation for its decision.
@@
 | Explainable | Explainable | ||||||||
-| F | Acc. | F | Acc. | F | Acc. | F | Acc. | ||
+| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | ||
 | Zero-Shot | No Definition | 20.48 | 52.70 | 61.43 | 53.28 | 64.96 | 58.40 | 63.75 | 60.41 |
@@
 | 8 (mixed domain) | 56.93 | 63.49 | 64.60 | 62.99 | 63.33 | 64.37 | 63.87 | 63.76 | |
+4 Evaluating Model Reliability in Detecting Framing
+The evaluation results for the GPT models across different settings are presented in Table 1, while the results for the FLAN-T5 models can be found in Appendix A.
+We can summarise the notable findings of the results as follows:
+Explainable Prompting Enhances Reliability:
+Explainable prompting consistently yielded more reliable outcomes in both zero- and few-shot variations, as shown by the reduced variance in accuracy and F1 scores across different prompting settings.
+Optimal Performance in Few-Shot with Diverse Examples and Explanation:
+The GPT-4 model achieved the highest accuracy and F1 scores in few-shot scenarios with a wide range of in-domain examples (8 varied in-domain).
+This result suggests that when the topic of news headlines is known in advance, incorporating this information into few-shot examples can considerably improve performance.
+Challenges with Cross-Domain Examples:
+F1 scores dropped in cross-domain settings without explainable prompts, falling below those in zero-shot settings with definitions.
+This highlights the challenges of applying advanced models to new or mixed domains, where headline topics vary widely.
+Therefore, for new domain applications, we suggest to either employ a zero-shot setting with task definition, or opting for explainable prompts when examples are used in the input prompts.
+We test our suggested settings on real world conditions in our analysis in section 5.
+Inherent Challenge in Framing Detection:
+Across both models, the highest F1 scores observed were in the range of 60-70%, highlighting the inherent difficulty of detecting framing.
+Framing detection requires understanding subtle language nuances, contextual cues, and the intended message framing, which are complex and context-dependent tasks.
+Furthermore, the performance between GPT-3.5 and GPT-4 did not significantly differ, particularly in more realistic use cases like zero-shot and cross-domain settings.
+Considering the higher cost of using GPT-4’s API, GPT-3.5 might be a more cost-effective choice for framing detection in large scale analysis.
+FLAN-T5 Performance Lag:
+Out-of-the-box FLAN-T5 models significantly underperform compared to GPT models, with the best F1 score reaching only 51%, which is negligibly above the majority baseline.
+The detailed results can be found in Appendix A.
+This performance lag indicates that smaller models like FLAN-T5 may not yet be viable for complex tasks such as framing detection without additional task-specific fine-tuning or more extensive training data.
+To assess this, we have fine-tuned the best performing model (FLAN-T5 Large) on the GVFC and tested it.
+This analysis is discussed in section 5.
+5 Analysis
+In this section, we perform a set of additional analysis and experiments to identify potential biases and areas for future research enhancements.
+The Impact of Emotional Language on Framing Detection
+Our analysis of GPT-4’s errors reveals a consistent pattern: the model frequently interprets emotional language as an indicator of framing, irrespective of whether the headline is actually framed or not.
+This tendency likely stems from the observation that emotionally charged language often accompanies framed headlines, evoking specific frames and eliciting emotional responses from readers, thereby influencing their perception of the news narrative Valkenburg et al.
+(1999).
+However, this leads to a biased error in the model, mistakenly identifying expressions of emotion as indicative of framing.
+Our findings are consistent with recent research on the weight of emotional charge in sentiment classification tasks with LLMs Burnham (2024).
+Such a pattern highlights the challenge of distinguishing between genuine emotional expression and intentional framing.
+Making this distinction is essential for enhancing the model’s precision in analysing news narratives.
+Appendix C provides examples of such misclassifications by GPT-4 in the explainable zero-shot setting.
+On the other hand, GPT-3.5 seems to be less affected by emotional language.
+| Flan-T5 | GPT-3.5 | GPT-4 | ||||||||
+| Small | Base | Large | ||||||||
+| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | |
+| Clear | 2.70 | 44.62 | 10.39 | 46.92 | 51.61 | 42.31 | 71.43 | 70.15 | 75.81 | 77.61 |
+| Contested | 2.33 | 36.36 | 0 | 34.85 | 67.47 | 59.09 | 41.43 | 38.81 | 3.60 | 20.15 |
+Clear vs. Contested Cases of Framing Annotations
+During our initial analysis, we observed discrepancies in the GVFC data annotations, with some annotations being potentially incorrect.
+Additionally, there were ambiguous cases where multiple labels could be justified based on different interpretations.
+In this regard, we manually examined half of the headlines in GVFC, i.e., 1300 headlines.
+Within this subset, we identified 134 contested annotations.
+Examples include “Live: Trump visits Pittsburgh after synagogue shooting” and “Shopify bans sale of certain firearms, accessories”, both annotated as framed.
+However, not all 134 cases are necessarily wrong annotations.
+Some of them, such as “Thousands gather to honor victims of the mass shooting with tears, candlelight, and song”, present more nuanced challenges in framing detection.
+In this dataset, 63.4% of headlines in this subset are annotated as framed.
+Additionally, we selected another set of 134 headlines where the presence or absence of framing was more apparent.
+For instance, headlines such as “Two dead including shooter at Florida yoga studio” and “Parkland school shooter blames massacre on a ’demon’ voice” serve as clear examples of non-framed and framed headlines that are also annotated correctly, respectively.
+We call this subset the “Clear” subset.
+Table 2 reports the results of evaluating our models on these two distinct subsets.
+As expected, both GPT models exhibited strong performance on the Clear subset.
+However, they have a considerably low results on the contested subset, with GPT-4’s F1 score dropping to 3.60%.
+We have also calculated how often GPT-4 and GPT-3.5’s predictions aligned within these subsets.
+These agreement ratios are shown in Table 3.
+We observe that while the models reached an agreement of about 59% on the contested subset, only 16% of these agreed-upon predictions matched the gold labels.
+Having manually checked the annotations for the contested subset, our finding suggests that these models’ could be useful in flagging potential annotation inaccuracies within existing or new datasets, primarily when there is agreement between them.
+As an example, both models identify the “ Muslim Americans raise more than $200,000 for those affected by Pittsburgh synagogue shooting” as not framed while it is annotated as framed in the data.
+| Clear | Contested | |
+| Agreement | 69.40 | 58.95 |
+| Agreement GL | 58.20 | 16.45 |
+| GPT-3.5 Turbo | GPT-4 | |||||||
+| Explainable | Explainable | |||||||
+| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | |
+| No Definition | 20.48 | 52.70 | 61.43 | 53.28 | 64.96 | 58.40 | 63.75 | 60.41 |
+| Different Wording | 25.39 | 53.55 | 59.04 | 54.64 | 47.40 | 60.56 | 45.55 | 57.76 |
+| +Definition | 51.55 | 59.14 | 59.03 | 58.79 | 65.36 | 63.84 | 64.64 | 60.99 |
+| Different Wording | 58.64 | 58.16 | 63.56 | 59.44 | 57.03 | 62.63 | 59.02 | 61.07 |
+Impact of Prompt Variability
+One of the challenges with prompt-based techniques is their sensitivity to the wording of input prompts; even minor variations can lead to drastically different performances.
+We examined this variability by using various input examples in focused in-domain, varied in-domain, cross-domain, and mixed-domain settings as shown in Table 1.
+To further examine the impact of prompt wording on framing detection results, we altered the prompt from “Decide whether this claim is framed” to “Decide whether this claim contains framing bias” in our zero-shot experiments.
+Table 4 shows these findings.
+The results show that even a simple change in the input prompt can result in a notable performance difference.
+For instance, there was a 15-point drop in F1 score for GPT-4 when the prompt did not include a definition of framing.
+However, the performance variance is considerably lower when the prompts include a definition of framing.
+Evaluation In The Wild (ITW) Dataset
+Until now, our results and analysis were centered on the GVFC dataset, which is limited to the narrow subject of gun violence within the U.S.
+To assess the models in a scenario closer to real-world conditions, where the topics of headlines are varied and unknown, we collected a set of framed headlines from a website known for its emphasis on news framing777https://newsframes.wordpress.com/category/headlines/.
+To make the dataset more balanced, we have also collected instances of not framed news headlines from newspapers such as BBC, The Guardian and Daily Mail.
+In this dataset, an example of framed headline is “EU plot to scrap Britain”, while an example of not framed headline is “Around 80 migrants including children rescued from Channel”.
+The entire dataset consists of 157 headlines spanning diverse subjects such as British weather, health, migration and European matters, and it contains 83 framed headlines and 74 not framed headlines, with a Majority Baseline Accuracy of 52.86%.
+The annotations were confirmed by one of the authors, who is an expert in framing.
+We assessed the models on this novel dataset employing both zero and few-shot settings.
+For the zero-shot, motivated by our results in the different wording evaluation (Table 4), we used both the standard and explainable settings of the Zero-Shot + Definition prompt.
+For the few-shot settings with 8 different examples, we used the examples employed in the varied in-domain and cross-domain settings of the Experiments in Section 4.
+Table 5 shows the results of our In the Wild evaluation, where GPT-4 reached 80.25% accuracy in one of the Few-Shot settings.
+However, GPT-4 remains sensitive to the examples provided in the prompt, as shown by the 6.6 points difference between the two Few-Shot settings.
+| GPT-3.5 Turbo | GPT-4 | |||
+| F1 | Acc. | F1 | Acc. | |
+| ZS + definition (S) | 74.16 | 70.70 | 78.61 | 72.61 |
+| ZS + definition (E) | 74.15 | 66.24 | 78.00 | 71.97 |
+| FS8 GV-domain (E) | 73.83 | 75.16 | 77.84 | 73.89 |
+| FS8 cross-domain (E) | 75.68 | 71.34 | 82.49 | 80.25 |
+Fine-tuned FLAN-T5
+Using the GVFC we have trained FLAN-T5 for framing detection888To fine-tune FLAN-T5, the GVFC was split in train (70%) - dev (10%) - test (20%).
+Hyperparameters are described in Appendix B.
+It is important to note that in the experiments of Table 1 we use the entire GVFC dataset as the test set, whereas in this experiment the GVFC test set is considerably smaller to enable fine-tuning.: after task-specific fine-tuning, FLAN-T5 Large has reached an accuracy of 78.83% on the GVFC test set, performing better than the GPT models, as shown in table 6.
+The table also shows the results of FLAN-T5 Large’s performance on the ITW dataset, compared to the out-of-the-box GPT models: while fine-tuned FLAN-T5 outperforms both GPT models on the GVFC test set, it falls behind on a more varied dataset that does not contain headlines related to the issue of Gun Violence, reaching an accuracy of 68.78% compared to GPT-4’s accuracy of 80.25%.
+| GPT-3.5-Turbo | GPT-4 | FLAN-T5 Large | |
+| fine-tuned | |||
+| Acc. | Acc. | Acc. | |
+| GVFC | 65.06 | 67.18 | 78.73 |
+| ITW | 71.34 | 80.25 | 68.78 |
+6 Conclusions
+This work advances our understanding of the performance of pre-trained models, including GPT-4, GPT-3.5 Turbo, and Flan-T5 (small, base, and large), in detecting framing within news headlines without the need for additional fine-tuning, given the scarcity of annotated data for this task.
+Through a comprehensive experimental approach we found that GPT-4 excels in few-shot scenarios and when asked to explain its results.
+However, it also showed a tendency to classify headlines as framed when they contain emotional language, while GPT-3.5 showed more resilience against such biases.
+Our results indicate that the models are sensitive to the wording of the prompt and the selected examples.
+However, including a linguistic definition of framing and prompting the model to explain its results leads to more reliable performances.
+We verify this finding by collecting a new dataset that better resembles real-world use cases for framing detection, where the test data encompasses various topics and the topics are not known in advance.
+Furthermore, our study also identifies a performance shortfall in Flan-T5, underscoring the challenges smaller models face in complex detection tasks without specific fine-tuning.
+As we have shown, after fine-tuning, FLAN-T5’s performance increases surpassing GPT models on in-domain data.
+However, the situation is reversed again when facing more diverse data (ITW), suggesting that, when the domain is not known in advance, fine-tuning might not be the best choice.
+Finally, through our analysis of the GVFC annotations and clear/contested subsets, we noticed the potential of these models in flagging annotation inaccuracies in new or existing datasets, when the predictions are consistent across multiple models.
+This would be particularly effective when checking annotations, if there is strong agreement between models on a label that contrasts the gold label.
+7 Limitations
+The findings of this study have to be seen in light of some limitations.
+For example, our evaluation focuses solely on English-language content, leaving space for further investigation on other languages to explore our findings’ applicability to non-English contexts.
+This limitation suggests a need for further investigation into the performance of LLMs across different languages and cultural contexts to fully assess the potential use of these models in social science research for detecting framing and analysing media narratives.
+Furthermore, two of the five models evaluated in our work are accessible only through OpenAI’s API, which is closed-source and subject to changes over time.
+This could affect the reproducibility of our results with newer versions of API, and they may have their own limitations.
+Therefore, focusing on improving open-source models emerges as a critical pathway forward, ensuring broader accessibility and reproducibility in research.
+Appendix
+Appendix A FLAN-T5 Performance Lag
+Table 7 shows the results of FLAN-T5 without any specific fine-tuning.
 | FLAN-T5 Small | FLAN-T5 Base | FLAN-T5 Large | |||||||||||
 | Explainable | Explainable | Explainable | |||||||||||
-| F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | ||
-| Zero-Shot | No Definition | 0.46 | 49.06 | 17.96 | 49.21 | 0 | 0 | 25.86 | 52.74 | 34.33 | 46.19 | 51.01 | 49.06 |
-| +Definition | 0.15 | 50.13 | 1.21 | 50.10 | 0 | 0 | 0.61 | 50.06 | 42.27 | 48.06 | 37.63 | 43.96 | |
-| Few-Shot | 2 examples | 0.31 | 50.06 | 0.31 | 50.06 | 0 | 0 | 3.12 | 49.98 | 42.27 | 48.06 | 37.63 | 43.96 |
-| 8 (focused in-domain) | 0.15 | 50.17 | 3.84 | 50.10 | 0 | 0 | 4.83 | 50.17 | 16.78 | 47.53 | 36.85 | 44.69 | |
-| 8 (varied in-domain) | 0 | 0 | 0.15 | 49.94 | 0 | 0 | 1.66 | 49.94 | 22.78 | 45.96 | 35.03 | 45.27 | |
-| 8 (cross domain) | 5.30 | 49.37 | 2.95 | 49.60 | 0 | 0 | 4.99 | 50.33 | 40.23 | 45.11 | 51.64 | 46.88 | |
-| 8 (mixed domain) | 0.31 | 50.21 | 1.51 | 50.02 | 0 | 0 | 3.84 | 50.13 | 34.39 | 44.58 | 44.05 | 43.43 | |
-4 LLMs’ Reliability in Detecting Framing Bias
-The evaluation results for GPT and FLAN-T5 models across different settings are presented in Table 1 and Table 2, respectively. We can summarise the notable findings of the results of as follows:
-Explainable Prompting Enhances Reliability:
-Explainable prompting consistently yielded more reliable outcomes in both zero- and few-shot variations, as shown by the reduced variance in accuracy and F scores.
-This underscores the importance of explainable settings for social science research on framing bias.
-Optimal Performance in Few-Shot with Diverse Examples:
-The GPT-4 model achieved the highest accuracy and F scores in few-shot scenarios with a wide range of in-domain examples (8 varied in-domain).
-This finding suggests that incorporating diverse, relevant in-domain examples significantly improves model performance when the news headline domain is known.
-Tendency Towards Framing Bias Classification:
-In explainable settings, both GPT-3.5 and GPT-4 models exhibit a higher tendency to classify headlines as framed.
-This is indicated by higher F scores for framed headline detection alongside lower overall accuracy, suggesting a bias towards classifying headlines as framed when prompted to identify framing.
-Challenges with Cross-Domain Examples:
-F scores dropped in cross-domain settings without explainable prompts, falling below those in zero-shot settings with definitions.
-This highlights the challenges of applying advanced models to new or mixed domains, where headline topics vary widely.
-Therefore, for new domain applications, we suggest to either employ a zero-shot setting with task definition, or opting for explainable prompts when examples are used in the input prompts.
-However, for large-scale data analysis, opting for a zero-shot approach without explanations could be more efficient, albeit at the potential cost of reduced reliability compared to the explainable setting.
-Inherent Challenge in Framing Bias Detection:
-Across both models, the highest F scores observed were in the range of 60-70 points, highlighting the inherent difficulty of detecting framing bias.
-Framing detection requires understanding subtle language nuances, contextual cues, and the intended message framing, which are complex and context-dependent tasks.
-Furthermore, the performance between GPT-3.5 and GPT-4 did not significantly differ, particularly in more realistic use cases like zero-shot and cross-domain settings.
-Considering the higher cost of using GPT-4’s API, GPT-3.5 might be a more cost-effective choice for framing bias detection in large scale analysis.
-FLAN-T5 Performance Lag:
-The results from Table 2 show that FLAN-T5 models significantly underperform compared to GPT models, with the best F score reaching only 51 points.
-This indicates that smaller models like FLAN-T5 may not yet be viable for complex tasks such as framing bias detection without additional task-specific fine-tuning or more extensive training data.
-5 Analysis
-In this section, we perform a set of additional analysis and experiments to identify potential biases and areas for future research enhancements.
+| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | ||
+| ZS | No Definition | 0.46 | 49.79 | 17.96 | 49.21 | 0 | 50.10 | 25.86 | 52.74 | 34.33 | 46.19 | 51.01 | 49.06 |
+| +Definition | 0.15 | 50.13 | 1.21 | 50.10 | 0 | 50.10 | 0.61 | 50.06 | 42.27 | 48.06 | 37.63 | 43.96 | |
+| FS | 2 examples | 0.31 | 50.06 | 0.31 | 50.06 | 0 | 50.10 | 3.12 | 49.98 | 5.74 | 49.64 | 8.32 | 49.64 |
+| 8 (focused in-domain) | 0.15 | 50.17 | 3.84 | 50.10 | 0 | 50.13 | 4.83 | 50.17 | 16.78 | 47.53 | 36.85 | 44.69 | |
+| 8 (varied in-domain) | 0 | 50.13 | 0.15 | 49.94 | 0 | 50.13 | 1.66 | 49.94 | 22.78 | 45.96 | 35.03 | 45.27 | |
+| 8 (cross domain) | 5.30 | 49.37 | 2.95 | 49.60 | 0 | 50.10 | 4.99 | 50.33 | 40.23 | 45.11 | 51.64 | 46.88 | |
+| 8 (mixed domain) | 0.31 | 50.21 | 1.51 | 50.02 | 0 | 50.10 | 3.84 | 50.13 | 34.39 | 44.58 | 44.05 | 43.43 | |
+Appendix B Detailed Evaluation Fine-tuned FLAN-T5 and Hyperparameters
+Table 8 presents the detailed performance comparison on the GVFC Test split and on the ITW dataset; the models are out-of-the-box GPT-3.5-Turbo and GPT-4, and fine-tuned FLAN-T5 Large.
+Table 9 presents the detailed performance comparison of fine-tuned FLAN-T5 small, base and large on the GVFC test set.
+The FLAN-T5 models (small, base, and large) were fine-tuned using the following hyperparameters: All models employ a learning rate of 5e-05, with a maximum input length of 70 tokens, a maximum label length of 4 tokens, and a batch size of 64.
+The number of epochs varies with 10 for the small, 40 for the base, and 17 for the large model, each within a 50-epoch limit and with early stopping set at 5 epochs to prevent overfitting.
+We used one A100 80G GPU, requiring less than 15 minutes of GPU time per model.
+| GPT-3.5-Turbo | GPT-4 | FLAN-T5 Large | ||||||||||
+| fine-tuned | ||||||||||||
+| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | |
+| GVFC | 67.39 | 65.06 | 60.52 | 76.02 | 69.08 | 67.18 | 62.42 | 77.33 | 76.33 | 78.83 | 80.63 | 72.46 |
+| ITW | 75.68 | 71.34 | 68.63 | 84.34 | 82.49 | 80.25 | 77.66 | 87.95 | 69.18 | 68.78 | 72.36 | 66.26 |
+| FLAN-T5 Small | FLAN-T5 Base | FLAN-T5 Large | ||||||||||
+| fine-tuned | fine-tuned | fine-tuned | ||||||||||
+| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | |
+| GVFC | 0 | 52.29 | 0 | 0 | 79.77 | 79.50 | 74.82 | 85.42 | 76.33 | 78.83 | 80.63 | 72.46 |
+| ITW | 2.38 | 47.70 | 100 | 1.20 | 63.87 | 56.05 | 56.48 | 73.49 | 69.18 | 68.78 | 72.36 | 66.26 |
+Appendix C Impact of Emotional Language on Framing Detection
+Table 10 shows some examples of the impact of emotional language on framing detection.
 | Headline: This list of lives lost in mass shootings grows longer. Here are their names. |
@@
 | It does not provide any particular perspective or agenda. |
-| Headline: Pregnant woman who killed intruder in justified shooting now faces felony gun charge due to previous marijuana conviction. |
+| Headline: Devastated’ Tamera Mowry Confirms Niece Alaina Housley Died In Thousand Oaks Shooting |
 | Gold: Not Framed, GPT-4: Framed, GPT-3.5: Not Framed |
-| GPT-4 Explanation: The claim is framed. It selectively emphasizes on the individual’s status as a pregnant woman and the nature of the shooting |
-| as justified, while sidelining the legal issue of her previous marijuana conviction which is directly linked to the current felony charge. |
-| The framing encourages empathy for the person involved by prominently highlighting her pregnancy and the self-defense aspect, provoking an emotional response. |
-| GPT-3.5 Explanation: This claim appears to accurately represent the situation. It states that a pregnant woman killed an intruder in a justified |
-| shooting and is now facing a felony gun charge due to a previous marijuana conviction. The claim does not seem to contain any misleading or false information. |
+| GPT-4 Explanation: The claim is framed as it emphasizes the emotional state of Tamera Mowry (being ’devastated’) while confirming the death of her niece |
+| in the shooting. This can elicit an emotional response from the audience and influence their perception of the issue. |
+| GPT-3.5 Explanation: The claim is not framed because it is simply reporting the confirmation of Tamera Mowry about the death of her niece in the Thousand |
+| Oaks shooting.It does not select any specific aspects or favor a certain interpretation of the issue. |
 | Headline: Gunman upset by divorce is found dead after manhunt. |
@@
 | There is no apparent bias or manipulation in the framing of this claim. |
-The Impact of Emotional Language on Framing Bias Detection
-Our analysis of GPT-4’s errors reveals a consistent pattern: the model frequently interprets emotional language as an indicator of framing bias, irrespective of the headline’s actual framing.
-This tendency likely stems from the observation that emotionally charged language often accompanies framed headlines, evoking specific frames and eliciting emotional responses from readers, thereby influencing their perception of the news narrative Valkenburg et al.
-(1999).
-However, this leads to a biased error in the model, mistakenly identifying expressions of emotion as indicative of framing bias.
-Such a pattern highlights the challenge of distinguishing between genuine emotional expression and intentional framing bias.
-Making this distinction is essential for enhancing the model’s precision in analysing news narratives.
-Table 3 provides examples of such misclassifications by GPT-4 in the explainable zero-shot setting.
-On the other hand, GPT-3.5 seems to be less affected by the emotional language.
-| GPT-3.5 | GPT-4 | Flan-T5 | ||||||||
-| Small | Base | Large | ||||||||
-| F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | |
-| Clear | 71.43 | 70.15 | 75.81 | 77.61 | 3.23 | 42.86 | 9.23 | 43.81 | 49.6 | 40.00 |
-| Contested | 41.43 | 38.81 | 3.60 | 20.15 | 3.03 | 39.62 | 0 | 0 | 65.60 | 59.43 |
-| Clear | Contested | |
-| Agreement | 69.40 | 58.95 |
-| Agreement GL | 58.20 | 16.45 |
-| GPT-3.5 | GPT-4 | |||||||
-| Explainable | Explainable | |||||||
-| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | |
-| No Definition | 20.48 | 52.70 | 61.43 | 53.28 | 64.96 | 58.40 | 63.75 | 60.41 |
-| Different Wording | 25.39 | 53.55 | 59.04 | 54.64 | 47.40 | 60.56 | 45.55 | 57.76 |
-| +Definition | 51.55 | 59.14 | 59.03 | 58.79 | 65.36 | 63.84 | 64.64 | 60.99 |
-| Different Wording | 58.64 | 58.16 | 63.56 | 59.44 | 57.03 | 62.63 | 59.02 | 61.07 |
-| Framed | Not-Framed | |
-| GPT-3.5 Turbo | 84 | 38 |
-| GPT-4 | 108 | 22 |
-| Flan-T5 Small | 1 | 129 |
-| Flan-T5 Base | 0 | 130 |
-| Flan-T5 Large | 32 | 98 |
-Clear vs. Contested Cases of Framing Annotations
-During our initial analysis, we observed discrepancies in some data annotations, with some annotations being potentially incorrect.
-Additionally, there were ambiguous cases where multiple labels could be justified based on different interpretations.
-In this regard, we manually examined half of the headlines in GVFC, i.e., 1300 headlines.
-Within this subset, we identified 134 contested annotations.
-Examples include “Live: Trump visits Pittsburgh after synagogue shooting” and “Shopify bans sale of certain firearms, accessories”, both annotated as framed.
-However, not all 134 cases are necessarily wrong annotations.
-Some of them, such as “Thousands gather to honor victims of the mass shooting with tears, candlelight, and song”, present more nuanced challenges in framing detection.88863.4% of headlines in this subset are annotated as framed in the dataset.
-Additionally, we selected another set of 134 headlines where the presence or absence of framing bias was more apparent.
-For instance, headlines such as “Two dead including shooter at Florida yoga studio” and “Parkland school shooter blames massacre on a ’demon’ voice” serve as clear examples of non-framed and framed headlines that are also annotated correctly, respectively.
-We call this subset the “Clear” subset.
-Table 4 reports the results of evaluating our models on these two distinct subsets.
-As expected, both GPT models exhibited strong performance on the Clear subset.
-However, they have a considerably low results on the contested subset, with GPT-4’s F score dropping to 3.6 points.
-We have also calculated how often GPT-4 and GPT-3.5’s predictions aligned within these subsets.
-These agreement ratios are shown in Table 5.
-We observe that while the models reached an agreement of about 59% on the contested subset, only 16% of these agreed-upon predictions matched the gold labels.
-This suggests that these models’ can be useful in flagging potential annotation inaccuracies within existing or new datasets.
-As an example, both models identify the “ Muslim Americans raise more than $200,000 for those affected by Pittsburgh synagogue shooting” as not framed while it is annotated as framed in the data.
-Impact of Prompt Variability
-One of the challenges with prompt-based techniques is their sensitivity to the wording of input prompts; even minor variations can lead to drastically different performances.
-We examined this variability by using various input examples in focused in-domain, varied in-domain, cross-domain, and mixed-domain settings as shown in Table 1.
-To further examine the impact of prompt wording on framing bias detection results, we altered the prompt from “Decide whether this claim is framed or not” to “Decide whether this claim contains framing bias” in our zero-shot experiments.
-Table 6 shows these findings.
-Our result indicates that the explainable setting yields the most consistent results, where different formulations of the question minimally affects outcome variability.
-On the other hand, we observe the most significant performance change in the GPT-4 model without a defined task or explanation, where the F score drops from 64 to 47 points.
-Given that the explainable setting requires more computational resources, including a task definition is a more efficient alternative for decreasing variability in results, and therefore enhancing their reliability.
-Evaluation in the Wild
-Until now, our results and analysis were centered on the GVFC dataset, which is limited to the narrow subject of gun violence within the U.S.
-To assess the models in a scenario closer to real-world conditions, where the topics of headlines are varied and unknown, we collected a new dataset from a website known for its emphasis on news framing.999https://newsframes.wordpress.com/category/headlines/ This collection consists of 130 headlines spanning diverse subjects such as British weather, health, and European matters.
-We assessed the models on this novel dataset employing the explainable few-shot setting with 8 cross-domain examples, which resulted in reliable results in Section 4.
-Table 7 shows the counts of headlines predicted as framed or not framed by each model.
-Given the dataset’s origin from a website known for featuring framed headlines, we anticipated the majority of headlines to be framed.
-We observe that the GPT-4 model predominantly identified headlines as framed.
-However, considering GPT-4’s tendency towards classifying inputs as framed, a promising avenue for future research involves developing a dataset specifically aimed at challenging the models with not-framed cases.
-This could include not-framed headlines that still incorporate emotional language, as discussed in the beginning of this section.
-6 Conclusions
-This work advances our understanding of the performance of LLMs such as GPT-3.5 Turbo, GPT-4, and Flan-T5 (small, base, and large) in detecting framing bias within news headlines.
-Through a comprehensive experimental approach that included zero-shot, few-shot, and explainable prompting settings, we found that GPT-4 excels in few-shot scenarios with a variety of in-domain examples, highlighting its superior capability in recognising framing bias.
-However, GPT-4 also displayed a tendency to classify headlines as framed, suggesting a valuable line of research could involve testing these models on datasets containing non-framed headlines that potentially also incorporate emotional language, identified as another potential weakness.
-In fact, our research explores the impact of emotional language on framing bias detection, with GPT-3.5 showing resilience against emotional language.
-The study also identifies a performance shortfall in Flan-T5, underscoring the challenges smaller models face in complex detection tasks without specific fine-tuning.
-Our findings indicate that F1 scores drop in cross-domain scenarios with non-explainable prompts, suggesting the effectiveness of zero-shot approaches with clear definitions or explainable prompts for new domain applications.
-The study further demonstrates the utility of these models in identifying potential annotation inaccuracies in new or already existing datasets.
-7 Limitations
-The findings of this study have to be seen in light of some limitations.
-For example, our evaluation focuses solely on English-language content, leaving space for further investigation on other languages to explore our findings’ applicability to non-English contexts.
-This limitation suggests a need for further investigation into the performance of LLMs across different languages and cultural contexts to fully assess the potential use of these models in social science research for detecting framing bias and analysing media narratives.
-Furthermore, two of the five models evaluated in our work are accessible only through OpenAI’s API, which is closed-source and subject to changes over time.
-This could affect the reproducibility of our results with newer versions of API, and they may have their own limitations.
-Therefore, focusing on improving open-source models emerges as a critical pathway forward, ensuring broader accessibility and reproducibility in research.
-8 Acknowledgements
-This work was supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1].
+Appendix D Detailed Results GPT Evaluation
+Tables 11 and 12 provide the detailed version of the results of the GPT Evaluation, including Precision and Recall.
+| GPT-3.5 Turbo | |||||||||
+| Explainable | |||||||||
+| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | ||
+| Zero-Shot | No Definition | 20.48 | 52.70 | 63.71 | 12.20 | 64.96 | 58.40 | 52.25 | 74.52 |
+| +Definition | 51.55 | 59.14 | 63.16 | 43.55 | 65.36 | 63.84 | 58.60 | 59.46 | |
+| Few-Shot | 2 examples | 26.83 | 54.16 | 66.06 | 16.83 | 58.25 | 65.84 | 59.32 | 73.98 |
+| 8 (focused in-domain) | 40.72 | 58.25 | 69.62 | 28.73 | 64.50 | 63.30 | 59.43 | 72.51 | |
+| 8 (varied in-domain) | 46.36 | 60.22 | 70.91 | 34.44 | 68.51 | 65.38 | 58.90 | 71.04 | |
+| 8 (cross domain) | 41.22 | 57.67 | 67.19 | 29.73 | 59.42 | 66.04 | 62.79 | 58.86 | |
+| 8 (mixed domain) | 56.93 | 63.49 | 69.25 | 48.34 | 63.33 | 64.37 | 61.82 | 67.64 | |
+| GPT-4 | |||||||||
+| Explainable | |||||||||
+| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | ||
+| Zero-Shot | No Definition | 64.96 | 58.40 | 56.05 | 77.22 | 63.75 | 60.41 | 58.71 | 69.73 |
+| +Definition | 65.36 | 63.84 | 62.63 | 68.34 | 64.64 | 60.99 | 59.03 | 71.43 | |
+| Few-Shot | 2 examples | 58.25 | 65.84 | 74.73 | 47.72 | 64.23 | 64.92 | 65.41 | 63.09 |
+| 8 (focused in-domain) | 64.50 | 63.30 | 62.36 | 66.80 | 66.44 | 63.96 | 60.06 | 71.51 | |
+| 8 (varied in-domain) | 68.51 | 65.38 | 62.75 | 75.44 | 70.41 | 66.92 | 63.61 | 78.84 | |
+| 8 (cross domain) | 59.42 | 66.04 | 73.63 | 49.81 | 59.09 | 64.61 | 69.86 | 51.20 | |
+| 8 (mixed domain) | 63.33 | 64.37 | 65.14 | 61.62 | 63.87 | 63.76 | 63.58 | 64.17 | |
+Appendix E Detailed Results In The Wild Evaluation
+Table 13 shows the detailed results of the evaluation In The Wild, including Precision and Recall.
+| GPT-3.5 Turbo | GPT-4 | |||||||
+| F1 | Acc. | Prec | Rec. | F1 | Acc. | Prec | Rec. | |
+| ZS + definition (S) | 74.16 | 70.70 | 69.47 | 79.52 | 78.61 | 72.61 | 66.95 | 95.18 |
+| ZS + definition (E) | 74.15 | 66.24 | 62.30 | 91.57 | 78.00 | 71.97 | 66.67 | 93.98 |
+| FS8 GV-domain | 73.83 | 75.16 | 83.33 | 66.27 | 77.84 | 73.84 | 70.59 | 86.75 |
+| FS8 cross-domain | 75.68 | 71.34 | 68.63 | 84.34 | 82.49 | 80.25 | 77.66 | 87.95 |